# Exploring the Impact of In-Browser Deep Learning Inference on Quality of User Experience and Performance

### v1

@@ Abstract.
Deep Learning (DL) is increasingly being integrated into Web applications through a method known as ”in-browser inference”, where the DL processes occur directly within Web browsers. However, the actual performance of this method and its effect on user experience quality (QoE) is not well-understood. This gap in knowledge necessitates new forms of QoE measurement, going beyond traditional metrics such as page load time. To address this, we conducted the first extensive performance evaluation of in-browser inference. We introduced new metrics for this purpose: responsiveness, smoothness, and inference accuracy.
Our thorough study included 9 widely-used DL models and tested them across 50 popular PC Web browsers. The findings show a significant latency issue with in-browser inference: it’s on average 16.9 times slower on CPU and 4.9 times slower on GPU than native inference methods. Several factors contribute to this latency, including underused hardware instruction sets, inherent delays in the runtime environment, resource competition within the browser, and inefficiencies in software libraries and GPU abstractions.
Moreover, in-browser inference demands a lot of memory, sometimes up to 334.6 times more than the size of the DL models themselves. This excessive memory usage is partly due to suboptimal memory management. Additionally, we noticed that in-browser inference increases the time it takes for graphical user interface (GUI) components to load in web browsers by a significant 67.2%, which severely impacts the overall QoE for users of web applications that depend on this technology.
copyright: nonecopyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXccs: Human-centered computing Empirical studies in ubiquitous and mobile computingccs: Computing methodologies Artificial intelligence
@@ 1. Introduction
Deep Learning (DL) has been increasingly adopted in various Web applications by deploying DL models in Web browsers for inference tasks, i.e., in-browser inference (Ma et al., 2019; web, 2023b; fac, 2023; aig, 2023; off, 2023; tea, 2023). In this context, DL frameworks and models are encapsulated and delivered within Web pages without additional installations, and DL inference performs directly in Web browsers (Ma et al., 2019).
The proliferation of DL-powered Web applications has made in-browser inference performance a critical topic for the Web research community. However, existing research efforts primarily focus on DL inference performance in native and server environment (ten, 2023; Liu et al., 2022a; Zhang et al., 2022; Chen et al., 2022; Gao et al., 2020). In contrast, there is still a notable knowledge gap about DL in-browser inference performance in the literature. In addition, such performance is essential for the quality of experience (QoE) during Web browsing. Moreover, no previous work has considered the impact of in-browser inference on the QoE. Existing Web QoE measurement efforts primarily focus on the loading phase of Web pages such as page load time (spe, 2023a; Bocchi et al., 2016; Tian and Ma, 2019; PLT, 2023; Ling et al., 2018; Yeo et al., 2019), while, essentially, the QoE of in-browser inference is related to the whole process of user’s browsing experience.
@@
To fill the knowledge gap, we present the first comprehensive measurement study of in-browser inference from two aspects: inference performance and QoE. Our extensive study involves 9 representative DL models across Web browsers of 50 popular PC devices. Furthermore, we use the most widely-adopted DL frameworks for deploying DL models in browsers, including TensorFlow.js (tfj, 2023b), ONNX Web Runtime (ort, 2023b).
@@
We observe a significant disparity in prediction latency between in-browser and native inference, averaging 16.9$\times$ on CPU and 30.6$\times$ on GPU. In the CPU case, this disparity arises primarily from the underutilization of Single-Instruction-Multiple-Data (SIMD) instructions, WebAssembly (Wasm) runtime-inherent inefficiency, as well as competition for resources with inherent browser tasks. On the GPU side, the inefficiency of software libraries on the Web, specifically WebGL (web, 2023a), compared to native CUDA (cud, 2023), contributes to the gap. Furthermore, our analysis reveals substantial performance variations across different devices, with disparities reaching up to 28.4$\times$ on CPU and 19.4$\times$ on GPU. These variations stem from the diverse capacities of devices and the hardware affinities of pre-designed DL kernels. Additionally, we identify ongoing concerns regarding memory footprint during in-browser inference. In certain scenarios, the memory footprint can escalate to 334.6$\times$ the model size during super-resolution model inference. These findings underscore the inefficiency of the in-browser inference on memory management when the intermediate data size increases.
@@ 2.1. In-browser Inference
Web browsers have emerged as a pivotal platform for the widespread adoption of DL technologies (Chen et al., 2020; Ma et al., 2019). DL models find applications across various Web applications, enabling in-browser inference tasks that encompass a range of capabilities such as grammar checking for online document (off, 2023), background blurring for virtual meetings (tea, 2023), extended reality (Bi et al., 2023), and enhancing online chat interactions (web, 2023b).
In-browser inference is the process where Web applications execute DL inference directly within Web browsers. Figure 1 illustrates the workflow of this process, which comprises three primary stages: setup, warmup, and prediction. During the setup stage, Web applications download the necessary DL models and DL frameworks and then load them into Web browsers. In the warmup stage, browsers, assisted by DL frameworks, allocate the required memory for inference and initiate the first DL inference operation. Finally, in the prediction stage, DL models carry out subsequent inference tasks using the capabilities provided by DL frameworks.
@@ 2.2. DL Frameworks and Backends
Both TF.js and ORT.js support using backends for inference acceleration, including WebAssembly (Wasm) (Haas et al., 2017; was, 2023c) on CPU and WebGL (web, 2023c) on GPU. Wasm serves as a binary instruction format for a stack-based virtual machine within Web browsers. It serves the purpose of a portable compilation target for programming languages, facilitating their deployment on the Web. Wasm stands out for its efficiency, owing to its compact size and swift loading capabilities (was, 2023c). In Wasm, Single-Instruction-Multiple-Data (SIMD) and multithreading are available for computation efficiency. SIMD is a hardware architecture that processes multiple data elements in one instruction, greatly enhancing computational speed. Multithreading is a technique that allows a program to execute multiple tasks concurrently, making efficient use of system resources for improving performance. WebGL is a JavaScript library designed for rendering graphics directly within Web browsers, implementing an API that adheres to the OpenGL standards (ope, 2023). Developers can harness the power of WebGL by crafting WebGL shaders (web, 2023a) to expedite parallel computation processes.
@@ 3.1. Experimental Setup
We start by outlining our choices for DL frameworks, backends, DL models, PC devices, and Web browser.
@@
Selection of PC devices and Web browser. Our measurement involves 50 different PC devices, featuring a diverse array of CPU models, including both Intel CPUs and Apple Silicon. Within the Intel CPU category, we encompass core-i5 and core-i7 processors spanning from Skylake to Rocket Lake architectures (sky, 2023; roc, 2023). In the Apple Silicon category, we consider M1 and M2 CPU models. Moreover, these devices are equipped with a range of GPU models, covering both integrated and discrete graphics solutions. The discrete GPU options encompass Nvidia models such as GTX 980, 1060, and 2060, while integrated GPUs are embedded within the CPU and include variants such as Intel HD 530, 630, and 645, among others. For the Web browser, we employ Google Chrome, which stands as the most widely used browser, commanding a global market share of more than 63.5% as of September 2023 (bro, 2023).
@@ 3.2. Measurement of Inference Performance
Latency. We measure the latency in both model-level and kernel-level. In terms of latency, the total inference latency consists of three types: setup latency, warmup latency, and prediction latency. (1) Model-level latency. We analyze the average latency gap and latency distribution among different DL frameworks, backends, and PC devices, and analyze the root cause. There are different acceleration techniques in the Wasm backend, i.e., SIMD and multithreading. The two DL frameworks that we adopt both support them for acceleration. SIMD processes multiple data within a single instruction and multithreading performs computation in parallel. To explore to what extent each acceleration technique reduces the latency, we measure the latency under different settings, i.e., whether each technique is enabled or not. (2) Kernel-level latency. A DL model can be represented as a series of interconnected kernels, such as matrix multiply and convolution. Understanding the performance of different kernels can help developers design kernel-specific optimization to improve the inference performance. For kernel-level latency, we measure the average kernel latency proportion on different backends. We do not consider the setup stage in this part because this stage does not involve kernel execution. In addition, we carry out a breakdown analysis to investigate how much SIMD and multithreading contribute to the acceleration in the Wasm backend.
@@ 3.3. Definition of QoE Metrics Dedicated to In-browser Inference
Responsiveness. Users prefer Web pages that respond to user requests quickly. The responsiveness reflects how quickly a Web page responds to user requests. It is quantified by the response time to user requests. We measure the responsiveness of Web pages based on Speedometer (spe, 2023b), a benchmark for Web application responsiveness. Speedometer simulates various Web activities, including 480 different tasks. As defined in Eq.1, we use the score “runs per minute” to quantify the responsiveness. The score represents the number of benchmark tests completed per unit time. We inject the inference code into the Speedometer Web page and perform in-browser inference during the benchmark test.
@@
Smoothness. Users prefer Web pages that render smoothly without any lag. The smoothness reflects how timely the GUI content is rendered. It is quantified by the rendering/video frames per second (fps). As defined in Eq.2, we measure the fps of playing video and rendering Web pages. The fps stands for the average number of rendered frames per unit time during the entire testing phase. For vision tasks, we select YouTube as the target Web page for measurement; for the text task, we select Google Docs. The two Websites are among the most popular Websites for online videos and documents. YouTube provides continuous video streams and we can input text continuously into Google Docs. We measure the video fps in YouTube and rendering fps in Google Docs. We inject the inference code into each Web page. We run the inference for 10 minutes while using the Web page and use the averaged results as the final obtained fps.
@@
Inference accuracy. Due to insufficient computation resource, inference accuracy might degrade inevitably. For example, video frames may be skipped and the application reuses previous results. Thus, as defined in Eq.3, we use inference accuracy, such as classification accuracy and detection mean average precision (mAP), to quantify the QoE. We use vision tasks, i.e., image classification and object detection, as the targets. We do not consider grammar checking in this part because, to our best knowledge, there lacks a benchmark dataset that contains both text and continuous timestamps. We use two benchmark datasets, including YouTube bounding boxes dataset (you, 2023a) for image classification and YouTube Face dataset (you, 2023b) for object detection. Both datasets contain continuous video frames and ground truth, as well as timestamps. Note that frames can use the previous inference results if they are not processed by DL models (Kang et al., 2017; Xu et al., 2020). We use the same models as before. We simulate in-browser inference with the fps obtained in the previous experiments.

### v2

# Anatomizing Deep Learning Inference in Web Browsers

@@ Abstract.
Web applications have increasingly adopted Deep Learning (DL) through in-browser inference, wherein DL inference performs directly within Web browsers. The actual performance of in-browser inference and its impacts on the quality of experience (QoE) remain unexplored, and urgently require new QoE measurements beyond traditional ones, e.g., mainly focusing on page load time. To bridge this gap, we make the first comprehensive performance measurement of in-browser inference to date. Our approach proposes new metrics to measure in-browser inference: responsiveness, smoothness, and inference accuracy. Our extensive analysis involves 9 representative DL models across Web browsers of 50 popular PC devices and 20 mobile devices. The results reveal that in-browser inference exhibits a substantial latency gap, averaging 16.9 times slower on CPU and 4.9 times slower on GPU compared to native inference on PC devices. The gap on mobile CPU and mobile GPU is 15.8 times and 7.8 times, respectively. Furthermore, we identify contributing factors to such latency gap, including underutilized hardware instruction sets, inherent overhead in the runtime environment, resource contention within the browser, and inefficiencies in software libraries and GPU abstractions. Additionally, in-browser inference imposes significant memory demands, at times exceeding 334.6 times the size of the DL models themselves, partly attributable to suboptimal memory management. We also observe that in-browser inference leads to a significant 67.2% increase in the time it takes for GUI components to render within Web browsers, significantly affecting the overall user QoE of Web applications reliant on this technology.
Deep learning, Web browser, measurement
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXccs: Human-centered computing Empirical studies in ubiquitous and mobile computingccs: Computing methodologies Artificial intelligence
@@ 1. Introduction
Deep Learning (DL) has been increasingly adopted in various Web applications by deploying DL models in Web browsers for inference tasks, i.e., in-browser inference (Ma et al., 2019; web, 2023b; fac, 2023; aig, 2023; off, 2024; tea, 2023). In this context, DL frameworks and models are encapsulated and delivered within Web pages without additional installations on user devices, and DL inference is performed directly in Web browsers (Ma et al., 2019), effectively streamlining user access to DL-powered functionalities. Many Web sites have already deployed in-browser inference. For example, online Office 365 (off, 2024) utilizes in-browser inference to assist with the online editing of documents and slides; online Teams blurs the background of people for video meetings (tea, 2023); Facetest (fac, 2023), Aige (aig, 2023), and Theremix (the, 2024) apply in-browser inference for tasks including image classification, face recognition, and music generation.
The proliferation of DL-powered Web applications makes in-browser inference a critical topic in various research communities, including Software Engineering (SE) (Chen et al., 2022; Gao et al., 2020; Guo et al., 2019; Cao et al., 2022; Liu et al., 2022a; Quan et al., 2022), Web (Ma et al., 2019; Zhang et al., 2022), and Mobile Computing (Tian et al., 2022; Huang et al., 2021). Firstly, the performance of in-browser inference is crucial to Web-based software applications. For instance, understanding the detailed performance of in-browser inference can assist software developers in refining the applications, thereby offering users a smoother and better service (Zhang et al., 2019; Yang et al., 2022). Moreover, the Quality of Experience (QoE) is pivotal in Web contexts (Upadhyaya et al., 2015), as it reflects the user’s experience while browsing Web pages. Improved QoE can attract a larger user base, making it vital for the development and deployment of software applications.
Existing research efforts primarily focus on DL inference performance in native [^1] We use “native” to refer to the non-browser environment, including servers, PCs, and mobile devices (Chen et al., 2022; Gao et al., 2020; Guo et al., 2019; Cao et al., 2022; Liu et al., 2022a; Zhang et al., 2022; Tang et al., 2023), regarding inference latency, memory footprint, etc. For instance, Tang et al. (Tang et al., 2023) accelerate DL model inference in PCs and mobile devices through a lookup table; Gao et al. (Gao et al., 2020) estimate the GPU memory footprint during DL model training in the server. However, in the scenario of in-browser inference, the browser acts as an intermediary, isolating Web applications from the underlying operating systems and restricting the application’s ability to fully harness hardware capabilities. This creates a distinction between the browser environment and the native environment. The current work demonstrates an insufficient exploration of in-browser inference and there are significant differences between browser and native environments, e.g., the absence of some advanced single instruction multiple data (SIMD) instructions. These factors contribute to a notable knowledge gap in the literature regarding in-browser inference performance. This gap underscores the need for more focused research to better understand and optimize the performance of in-browser inference. Moreover, existing Web QoE measurement efforts primarily focus on the loading phase of Web pages such as page load time (spe, 2023a; Bocchi et al., 2016; Tian and Ma, 2019; PLT, 2023; Ling et al., 2018; Yeo et al., 2019). For example, Speed Index (spe, 2023a) measures how quickly the Web page is visually displayed during page loading. To the best of our knowledge, no previous work has considered the impact of in-browser inference on the QoE.
@@
To fill the knowledge gap, we present the first comprehensive measurement study of in-browser inference from two aspects: inference performance and QoE. Our extensive study involves 9 representative DL models across Web browsers of 50 popular PC devices and 20 mobile devices. Furthermore, we use the most widely adopted DL frameworks for deploying DL models in browsers, including TensorFlow.js (tfj, 2023b), ONNX Web Runtime (ort, 2023b).
@@
We observe a significant disparity in prediction latency between in-browser and native inference, averaging 16.9$\times$ on CPU and 30.6$\times$ on GPU on PC devices, and the gap on mobile CPU and mobile GPU is 15.8$\times$ and 7.8$\times$, respectively. In the CPU case, this disparity arises primarily from the underutilization of SIMD instructions on PC devices, WebAssembly (Wasm) runtime-inherent inefficiency, as well as competition for resources with inherent browser tasks. On the GPU side, the inefficiency of software libraries on the Web, i.e., WebGL (web, 2023a), compared to native CUDA (cud, 2023), contributes to the gap. Furthermore, our analysis reveals substantial performance variations across different devices, with disparities reaching up to 28.4$\times$ on CPU and 19.4$\times$ on GPU. These variations stem from the diverse capacities of devices and the hardware affinities of pre-designed DL kernels. Additionally, we identify ongoing concerns regarding memory footprint during in-browser inference. In certain scenarios, the memory footprint can escalate to 334.6$\times$ the model size during super-resolution model inference. These findings underscore the inefficiency of the in-browser inference on memory management when the intermediate data size increases.
@@
To summarize, this paper makes the following contributions:
  • We conducted a detailed and granular analysis of in-browser inference performance at both the model and kernel levels across multiple devices, and examined the performance gap and underlying reasons compared to native inference.
  • We proposed specific Quality of Experience (QoE) metrics tailored for in-browser inference scenarios and conducted quantitative analyses to investigate the impact of in-browser inference on QoE.
  • We provided recommendations for browser vendors, inference framework developers, and Web application developers to help them enhance the performance of in-browser inference and improve user QoE.
  • As an additional contribution, we publicly released our code and data [^2] https://github.com/qipengwang/InBrowserInference to facilitate the replication and extension of this work.
@@ 2.1. In-browser Inference
Web browsers have emerged as a pivotal platform for the widespread adoption of DL technologies (Chen et al., 2020; Ma et al., 2019). Different from native applications that require installation and upgrade, raising barriers for end-users, and hindering AI service deployments, Web browsers, however, are promising platforms for DL adoption, without the requirement for extra installation. Besides, in-browser inference does not need to send the user data to the remote server, preserving user privacy. In-browser inference has attracted tremendous research attention (Chen et al., 2022; Gao et al., 2020; Guo et al., 2019; Cao et al., 2022; Liu et al., 2022a; Quan et al., 2022; Ma et al., 2019; Zhang et al., 2022; Tian et al., 2022; Huang et al., 2021; Jia et al., 2024; Dong et al., 2023). In-browser inference empowers many new abilities of Web applications. For example, Web Office 365 integrates grammar checking and typing completion (off, 2024); online Teams (tea, 2024) use in-browser inference to analyze facial contours for online meetings; Facetest (fac, 2023) applies in-browser inference for face detection; Aige (aig, 2023) applies in-browser inference to guess the age of users according to face image; Theremix (the, 2024) utilizes in-browser inference for music generation according to users’ control. Many existing Web applications can benefit from in-browser inference. For instance, online subtitle generation will be enabled for living applications; Web augmented reality, virtual reality, and mixed reality may be more vivid when leveraging DL.
In-browser inference is the process where Web applications execute DL inference directly within Web browsers. Figure 1 illustrates the workflow of this process, which comprises three primary stages: setup, warmup, and prediction. During the setup stage, Web applications download the necessary DL models and DL frameworks and then load them into Web browsers. In the warmup stage, browsers, assisted by DL frameworks, allocate the required memory for inference and initiate the first DL inference operation. Finally, in the prediction stage, DL models carry out subsequent inference tasks using the capabilities provided by DL frameworks. The warmup and prediction stages in the native environment are identical to those in the Web environment, while the setup stage differs between the two environments. In the native environment, the setup stage merely requires loading the model and framework from local storage, eliminating the need to download them from the server.
@@ 2.2. Frameworks and Backends
Both TF.js and ORT.js support using backends for inference acceleration, including WebAssembly (Wasm) (Haas et al., 2017; was, 2023c) on CPU and WebGL (web, 2023) on GPU. Wasm serves as a binary instruction format for a stack-based virtual machine (VM) within Web browsers. It serves the purpose of a portable compilation target for programming languages, facilitating their deployment on the Web. Wasm stands out for its efficiency, owing to its compact size and swift loading capabilities (was, 2023c). In Wasm, Single-Instruction-Multiple-Data (SIMD) and multithreading are available for computation efficiency. SIMD is a hardware architecture that processes multiple data elements in one instruction, greatly enhancing computational speed. Multithreading is a technique that allows a program to execute multiple tasks concurrently, making efficient use of system resources for improving performance. WebGL is a JavaScript library designed for rendering graphics directly within Web browsers, implementing an API that adheres to the OpenGL standards (ope, 2023). Developers can harness the power of WebGL by crafting WebGL shaders (web, 2023a) to expedite parallel computation processes.
@@ 3.1. Experimental Setup
We start by outlining our choices for DL frameworks, backends, DL models, PC devices, and Web browsers.
@@
Selection of devices and Web browsers. Our measurement mainly involves 50 different PC devices, featuring a diverse array of CPU models, including both Intel CPUs and Apple Silicon. Within the Intel CPU category, we encompass core-i5 and core-i7 processors spanning from Skylake to Rocket Lake architectures (sky, 2023; roc, 2023). In the Apple Silicon category, we consider M1 and M2 CPU models. Moreover, these devices are equipped with a range of GPU models, covering both integrated and discrete graphics solutions. The discrete GPU options encompass Nvidia models such as GTX 980, 1060, and 2060, while integrated GPUs are embedded within the CPU and include variants such as Intel HD 530, 630, and 645, among others. For the Web browser, we employ Google Chrome, which stands as the most widely used browser, commanding a global market share of more than 63.5% as of September 2023 (bro, 2023). To explore the difference between different browsers, we consider Firefox to be another Web browser platform. According to browser market share data (bro, 2023), Firefox holds the fourth largest share. We excluded Safari because it is only available on Apple devices, and we excluded Edge because it is also a Chromium-based browser, the same as Chrome. To explore the performance of in-browser inference on mobile devices, we involve 20 mobile devices. The mobile devices also feature a wide spectrum of SoC models, including both Snapdragon SoC models and Kirin SoC models. The Snapdragon SoC spans from Snapdragon 855 to Snapdragon 8 Gen 3; The Kirin SoC spans from Kirin 980 to Kirin 9000. On these mobile devices, we use Chrome as the experiment Web browser platform.
@@ 3.2. Measurement of Inference Performance
Experiment steps. In the experiment, we implemented the following procedures: (1) Kernel-level [^3] A DL model can be represented as a series of interconnected kernels, such as matrix multiply and convolution. The interconnections of these kernels form a computation graph. profiling enabling. We modified the inference framework to enable profiling kernel-level latency and memory footprint. (2) Web service deployment. We deployed a Web service that supports inference using various models through TF.js or ORT.js frameworks. (3) Model inference on devices. Devices accessed the service to perform model inference, during which we recorded the latency of each model inference. (4) We repeated step (3) multiple times and calculated the average latency and memory footprint on each device and browser, as well as the latency variance.
Latency. We measure the latency in both model-level and kernel-level. In terms of latency, the total inference latency consists of three types: setup latency, warmup latency, and prediction latency. (1) Model-level latency. We analyze the average latency gap and latency distribution among different DL frameworks, backends, and PC devices, and analyze the root cause. There are different acceleration techniques in the Wasm backend, i.e., SIMD and multithreading. The two DL frameworks that we adopt both support them for acceleration. SIMD processes multiple data within a single instruction and multithreading performs computation in parallel. To explore to what extent each acceleration technique reduces the latency, we measure the latency under different settings, i.e., whether each technique is enabled or not. (2) Kernel-level latency. Understanding the performance of different kernels can help developers design kernel-specific optimization to improve the inference performance. For kernel-level latency, we measure the average kernel latency proportion on different backends. We do not consider the setup stage in this part because this stage does not involve kernel execution. In addition, we carry out a breakdown analysis to investigate how much SIMD and multithreading contribute to the acceleration in the Wasm backend.
@@ 3.3. Definition of QoE Metrics Dedicated to In-browser Inference
Responsiveness. Users prefer Web pages that respond to user requests quickly. The responsiveness reflects how quickly a Web page responds to user requests. It is quantified by the response time to user requests. We measure the responsiveness of Web pages based on Speedometer (spe, 2023b), a benchmark for Web application responsiveness. Given the diverse application scenarios in the Web environment, it is difficult to have metrics generalize to all models or use cases. To address this challenge, the WebKit team officially provides the Speedometer dataset that covers various use cases as a benchmark. Specifically, this benchmark encompasses multiple high-level user journeys, including editing rich text, rendering charts, and reading news sites, with each journey further comprising multiple workloads for testing. In this benchmark, “runs per minute” is suggested to reflect the responsiveness QoE. As defined in Eq.1, we use the score “runs per minute” to quantify the responsiveness. The score represents the number of benchmark tests completed per unit time. We inject the inference code into the Speedometer Web page through the “Selenium” library, a library for automated Web application development and testing, allowing the developer to execute any JavaScript programs on any Web page. We perform in-browser inference during the benchmark test intermittently, to avoid the benchmark test being blocked if the model inference is performed continually. During the test, model inference accounts for 50% of the total duration.
@@
Smoothness. Users prefer Web pages that render smoothly without any lag. The smoothness reflects how timely the GUI content is rendered. It is quantified by the rendering/video frames per second (fps). As defined in Eq.2, we measure the fps of playing video and rendering Web pages. The fps stands for the average number of rendered frames per unit time during the entire testing phase. For vision tasks, we select YouTube as the target Web page for measurement; for the text task, we select Google Docs. The two Websites are among the most popular Websites for online videos and documents. YouTube provides continuous video streams and we can input text continuously into Google Docs. We measure the video fps in YouTube and render fps in Google Docs. We used the first recommended playback page on the YouTube homepage. Regarding Google Docs experiments, we use different input texts on the devices at varying speeds ranging from 20 to 60 words per minute. We inject the inference code into each Web page. We run the inference for 10 minutes while using the Web page and use the averaged results as the final obtained fps.
@@
Inference accuracy. Due to insufficient computation resource, inference accuracy might degrade inevitably. For example, video frames may be skipped and the application reuses previous results. Thus, as defined in Eq.3, we use inference accuracy, such as classification accuracy and detection mean average precision (mAP), to quantify the QoE. We use vision tasks, i.e., image classification and object detection, as the targets. We do not consider grammar checking in this part because, to our best knowledge, there lacks a benchmark dataset that contains both text and continuous timestamps. We use two benchmark datasets, including the YouTube bounding boxes dataset (you, 2023a) for image classification and the YouTube Face dataset (you, 2023b) for object detection. The two datasets are widely used (Fan et al., 2019; Liu et al., 2023; Cai et al., 2019; Wang et al., 2018a; Pfister et al., 2015; Schroff et al., 2015) because they both contain continuous real-world video stream frames and corresponding timestamps.
Experiment steps. In our experiment, we executed the following procedures: (1) Code injection using Selenium. We utilized the Selenium (sel, 2024) library to inject model inference code into various Web pages. (2) Model inference and QoE monitoring. After the Web pages were fully loaded, we used Selenium to initiate the model inference functions. The inference process was performed for a period, during which we recorded responsiveness and smoothness QoE. We enable multithreading and set 4 threads. (3) Continuous in-browser inference simulation. We simulated continuous in-browser inference using the two datasets and the fps values obtained from the smoothness experiment. If a video frame could be processed using the model, the inference results of that frame were used. If a frame could not be processed due to low fps or high resource usage, the results from the previous frame were reused. This approach is quite common in the field of video processing (Kang et al., 2017; Xu et al., 2020). We calculate the inference accuracy after all frames are processed.